### v1

# TempoFormer: A Transformer for Temporally-aware Representations in Change Detection

Abstract
Dynamic representation learning plays a pivotal role in understanding the evolution of linguistic content over time. On this front both context and time dynamics as well as their interplay are of prime importance. Current approaches model context via pre-trained representations, which are typically temporally agnostic. Previous work on modeling context and temporal dynamics has used recurrent methods, which are slow and prone to overfitting. Here we introduce TempoFormer, the fist task-agnostic transformer-based and temporally-aware model for dynamic representation learning. Our approach is jointly trained on inter and intra context dynamics and introduces a novel temporal variation of rotary positional embeddings. The architecture is flexible and can be used as the temporal representation foundation of other models or applied to different transformer-based architectures. We show new SOTA performance on three different real-time change detection tasks.
TempoFormer: A Transformer for Temporally-aware Representations in Change Detection
Talia Tseriotou1, Adam Tsakalidis1,2,3 Maria Liakata1,2 1Queen Mary University of London, 2The Alan Turing Institute, 3European Centre for the Development of Vocational Training t.tseriotou@qmul.ac.uk
1 Introduction
Linguistic data sequences are generated continuously over time in the form of social media posts, written conversations or documents that keep evolving (e.g. through regular updates). While a large body of work has been devoted to assessing textual units or sub-sequences in isolation – i.e. in emotion classification (Alhuzali and Ananiadou, 2021), ICD coding (Yuan et al., 2022), task-specific dialogue generation (Brown et al., 2024), irony and sarcasm detection Potamias et al. (2020) – such approaches leave significant historical (often timestamped) context unused. Fig. 1 provides an example from the task of identifying mood changes through users’ online content, where the last post in isolation cannot denote if there has been a ‘Switch’ in the user’s mood – the historical content provides important context for the user’s originally positive mood, enhancing the signal for a negative switch in their behaviour.
Figure 1: The historical content provides important context towards identifying the Switch in the user’s mood from positive to negative in the last post.
Dynamic representation learning approaches aim to tackle this challenge. Dynamic word embedding methods have been studied in the context of semantic change detection (Bamler and Mandt, 2017; Rosenfeld and Erk, 2018). While changes in this context occur over long time periods, dynamic representation learning has been explored in other more temporally fine-grained tasks such as event detection (Yan et al., 2019; Yang et al., 2019; Lai et al., 2020), fake news detection (Vaibhav et al., 2019; Raza and Ding, 2022; Kaliyar et al., 2021) and mental health condition detection (Sawhney et al., 2021b; Tsakalidis et al., 2022a; Tseriotou et al., 2023). Nevertheless, dynamic representation learning research remains largely task or even dataset specific.
Transformer-based injection. The above mentioned approaches have relied on either pre-trained contextualised representations or transformer-based model layers (Devlin et al., 2018; Liu et al., 2019) to fine-tune representations before feeding them into RNN and CNN-like architectures as they had been shown to outperform transformer-based models (Ji et al., 2021; Gao et al., 2021; Tsakalidis et al., 2022b). However, since LSTM-based systems tend to overfit small datasets, transformer-based methods that overcome this issue would be a preferable method (Yu et al., 2020). Yet so far adapting layers on top of a transformer fails to strike the right balance between representation learning and task dynamics (Li et al., 2022; Ng et al., 2023).
Temporal modelling. Although integration of time in language models has been explored for temporal adaption (Röttger and Pierrehumbert, 2021) in semantic change detection, (Rosin and Radinsky, 2022; Wang et al., 2023) there is not yet work that explores the abilities of transformers to model temporally distant textual sequences (streams). Recently LLMs have been shown to fall short in terms of temporal reasoning (Jain et al., 2023; Wallat et al., 2024), especially in event-event temporal reasoning (Chu et al., 2023). Here we make the following contributions:
  • $\bullet$
    We present a novel, temporally-aware BERT-based model (‘TempoFormer’) [^1] https://github.com/ttseriotou/tempoformer that models streams of chronologically ordered textual information accounting for their temporal distance. TempoFomer is the first such model to directly modify the transformer architecture, doing so in a flexible and task-agnostic manner.
  • $\bullet$
    We transform rotary position embeddings into rotary temporal embeddings that measure the temporal distance of sequential data points.
  • $\bullet$
    Contrary to prior work reliant on pre-trained contextual embeddings, we allow for adaptation of transformers towards the domain and the temporal aspects of a dataset. We show that TempoFormer can be used as the foundation in more complex architectures (e.g. involving recurrence), striking the right balance between modeling a post/utterance (context-aware) and the timeline-level dynamics. Moreover the TempoFormer upper layers are flexible and can be applied in different Transformer-based architectures.
  • $\bullet$
    We show SOTA performance on 3 change detection NLP tasks (longitudinal stance switch, identifying mood changes and identifying conversation derailment).
2 Related Work
Context-aware Sequential Models: Numerous social media related tasks such as rumour detection rely on chronologically ordered conversation threads Ma et al. (2020); Lin et al. (2021); Ma and Gao (2020). Moreover Ng et al. (2023) have shown lift in performance when using the full context of medical notes, rather than the discharge summary alone, for ICD coding. However context-aware sequential models have so far relied on recurrent networks or hierarchical attention (Li et al., 2020; Ma et al., 2020; Tsakalidis et al., 2022a) without exploring the dynamics between sentence level and stream level representations.
Longitudinal Modeling and Change Detection: In addition to the importance of the linguistic stream, longitudinal tasks rely on temporal dynamics to asses the progression and identify changes over time. In the case of (a) identifying changes in user mood (Tsakalidis et al., 2022b; Hills et al., 2023; Tseriotou et al., 2023) and suicidal ideation through social media (Sawhney et al., 2021a) change is relative to the temporal evolution of users’ mood over time and approaches have relied mostly on recurrence and on utterance-level pretrained language model (PLM) representations. Tseriotou et al. (2024) introduced a longitudinal variation of (b) stance detection Yang et al. (2022); Kumar and Carley (2019) for detecting shifts (changes) in the public’s opinion towards an online rumour. They used Sentence-BERT (Reimers and Gurevych, 2019) representations with integration of path signatures (Lyons, 1998) in recurrence. For (c) conversation topic derailment, previous work has relied on fine-tuning transformer-based models (Konigari et al., 2021), providing extended context in their input (Kementchedjhieva and Søgaard, 2021) or applying recurrence over the utterance (Zhang et al., 2019a) and context stream (Chang and Danescu-Niculescu-Mizil, 2019). In this work we integrate stream dynamics directly into the transformer and show the flexibility of our approach as the foundation of different longitudinal tasks.
Temporal Language Modeling: Many of the above tasks involve timestamps, which can enhance change detection through temporal dynamics. However, little research in NLP leverages time intervals and those who do assume equidistant time intervals between events (Ma and Gao, 2020; Tsakalidis and Liakata, 2020). Other work on temporal modelling has relied on hand crafted periodic task-specific time features (Kwon et al., 2013), concatenation of timestamp with linguistic representations (Tseriotou et al., 2023, 2024) or Hawkes temporal point process applied on top of recurrence (Guo et al., 2019; Hills et al., 2023). These approaches applied on top of LM representations miss the opportunity of training representations informed by temporal dynamics. Additionally, transformer-based models lack temporal sensitivity (Lazaridou et al., 2021; Loureiro et al., 2022). Rosin and Radinsky (2022) has conditioned attention weights on time, while Rosin et al. (2022); Wang et al. (2023) concatenated time tokens to text sequences. Although these methods create time-specific contextualised embeddings, they utilise absolute points in time rather than leveraging the temporal distance between units of textual information, important for context-aware and longitudinal tasks. Here we adapt the transformer attention mechanism to cater for the relative temporal aspect (§3.5).
Hierarchical Models: Long content modeling approaches have leveraged transformer or attention-based blocks hierarchically on long documents, on input chunks/sentences and then on the sequence of such chunks (Zhang et al., 2019c; Pappagari et al., 2019; Wu et al., 2021; Li et al., 2023). This produces chunk-level summary embeddings, which preserve both the local and global aspects of contextualised representations. Here we leverage such local and global context dynamics to more efficiently model linguistic streams.
3 Methodology
Here we introduce the TempoFormer architecture. We first provide the problem formulation (§3.1), followed by model overview (§3.2) and then discuss the various model components (§3.3-3.7).
3.1 Problem Formulation
A fundamental concept underpinning longitudinal tasks is that of timelines, $P$, defined as chronologically ordered units of information between two dates (Tsakalidis et al., 2022b), here either in the form of a sequence of users’ posts, a conversation or an online thread. Specifically the $c$-th timeline, $P^{c}$, consists of a series of posts [^2] We use terms posts and utterances interchangeably as the exact nature of the textual unit depends on the specific task., $u_{i}$, each with a corresponding timestamp, $t_{i}$. $P^{c}=[\{u_{0},t_{0}\},\{u_{1},t_{1}\},...,\{u_{N-1},t_{N-1}\}]$. The length of the timeline, $N$, can vary. We formulate the problem of assessing textual units in a timeline as early-stage, real-time classification, following Tseriotou et al. (2023). We map each timeline into $N$ training samples, that we call streams. Each stream contains a predefined window, $w$, of the most recent posts and a label for the most recent post: ([{$([\{u_{i-w+1},t_{i-w+1}\},...\{u_{i-1},t_{i-1}\},\{u_{i},t_{i}\}],l_{i})$.
3.2 TempoFormer Overview
Figure 2: TempoFormer Architecture on 5-post window.
Fig. 2 provides an overview of TempoFormer. Its hierarchical architecture consists of three main modules, temporally-aware enhancement in multi-head attention to model the temporal distance between posts and a classification head. The modules are: post-level (local) encoding3.3) – obtaining word-level representation of each post using BERT’s first 10 layers; stream (global) encoding3.4) – modeling the sequential and temporal interactions between posts; and context-enhanced encoding3.6) – fusing stream-awareness in post-level representations to make them context-aware.
3.3 Post-level Encoding (Local)
Each training instance is a stream consisting of the current post and its recent history, alongside corresponding timestamps: [$[\{\mathbf{u}_{i-w+1},t_{i-w+1}\},...\{\mathbf{u}_{i-1},t_{i-1}\},\{\mathbf{u}_{i},t_{i}\}]$, with a total of $w$ posts in a stream. Timestamps are ignored at this stage. This stream of posts is converted into a stream, $e$, of word-level embeddings of word sequence length $K$ via the word and position embedding layer of BERT: [$[\{\mathbf{e}_{1,i-w+1},\mathbf{e}_{2,i-w+1}...,\mathbf{e}_{K,i-w+1}\},...$$\{\mathbf{e}_{1,i},\mathbf{e}_{2,i}...,\mathbf{e}_{K,i}\}]$. Specifically, in this module, the posts of each stream pass without post-post interactions via the first 10 BERT layers, resulting in hidden word-level representations for each post. Note that since a post is part of multiple streams through their window, it will pass through the BERT layers as part of each corresponding stream. For each post $j$ (belonging to a stream $q$), the word-level representations from the $z$-th Transformer layer are denoted as: $\mathbf{H}^{z}_{j_{q}}$ Therefore at the 10-th layer we reconstruct the stream and form local stream representation: $[\mathbf{H}^{10}_{i-w+1},...,\mathbf{H}^{10}_{i}]$.
3.4 Stream-level Encoding (Global)
Inspired by Wu et al. (2021), who model long documents hierarchically by stacking transformer-based layers of sentence, document and document-aware embeddings, we build stream and context-enhanced layers on top of post-level representations. At the stream encoding layer, we capture inter-stream dynamics. Stream-level position embeddings (PE), $\mathbf{s}^{10}$, added after the 10-th layer, encode post order within the stream. By then passing the word-level stream PE representations to another BERT layer, we obtain word-level sequence-aware updated hidden representations $[\mathbf{H}^{11}_{1,i-w+1}$, …,$\mathbf{H}^{11}_{1,i}]$.
Next, we obtain the order-aware [CLS] token from the stream and apply Temporal Rotary Multi-head Attention (MHA), a proposed variation of RoFormer (Su et al., 2024), which accounts for the temporal rather than the sequential distance between posts (see §3.5). These context-aware, temporally-enhanced tokens are fed back to replace the respective [CLS] tokens in the hidden representations from the previous BERT layer, resulting in $[\mathbf{H}^{{}^{\prime}11}_{1,i-w+1}$, …,$\mathbf{H}^{{}^{\prime}11}_{1,i}]$. This enables the propagation of the learnt stream embeddings to the post-level.
3.5 Temporal Rotary Multi-Head Attention
BERT relies on positional embeddings to meaningfully encode the sequential order of words which are then fused via self-attention. Such embeddings are absolute (position-specific) and lack a relative sense. Su et al. (2024) proposed the Rotary Position Embeddings (RoPE) that incorporate the relative position between tokens within self-attention. Besides flexibility (in terms of sequence length generalisability), this introduces in the formulation intuitive inter-token dependency, which decays with increasing token distance. Given the attention formulation $\textrm{Attn}(\mathbf{Q},\mathbf{K},\mathbf{V})_{m}=\dfrac{\sum^{N}_{n=1}(\textrm{exp}(\mathbf{q}^{T}_{m}\mathbf{k}_{n}/\sqrt{d}))\mathbf{v}_{n}}{\sum^{N}_{n=1}(\textrm{exp}(\mathbf{q}^{T}_{m}\mathbf{k}_{n}/\sqrt{d}))}$, where $m$/$n$ denote the query/key positions, after applying RoPE self-attention, the $\mathbf{q}^{T}_{m}\mathbf{k}_{n}$ becomes:
[…]
where $R^{d}_{\theta,m}$ is the rotary matrix with $d$ embedding dimensions and the following formulation:
$R^{d}_{\theta,m}=\begin{pmatrix}[c]\textrm{cos}(m\theta_{1})&-\textrm{sin}(m\theta_{1})&0&0\\ \textrm{sin}(m\theta_{1})&\textrm{cos}(m\theta_{1})&0&0\\ \vdots&\ddots&\ddots&\vdots\\ 0&0&\textrm{cos}(m\theta_{d/2})&-\textrm{sin}(m\theta_{d/2})\\ 0&0&\textrm{sin}(m\theta_{d/2})&\textrm{cos}(m\theta_{d/2})\\ \end{pmatrix}$
where $\theta_{i}=10000^{-2(i-1)/d}i\in[1,2,...,d/2]$. The rotary matrix incorporates the relative position information through rotation of $q$ and $k$ based on their position in the sequence. The dot product decreases as the tokens move further apart. In Eq. 1, the formulation results in the relative position $(m-n)$, so the rotation between the 6-th and the 3-rd tokens is the same as between the 7-th and the 4-th ones.
Here, in order to model the temporal dynamics, we propose a novel variation of Eq. 1, named Temporal Rotary Multi-head Attention, making use of the relative position property. Instead of $R^{d}_{\theta,n-m}$, we reformulate the rotary matrix to model the temporal, rather than the positional differences, $R^{d}_{\theta,\mathbf{t_{n}-t_{m}}}$. We employ it on the stream-level using the [CLS] tokens to capture the stream global context through both the temporal and linguistic dynamics. The developed layer includes solely self-attention without the need for feed-forward and normalisation layers. In practice, since we measure time in seconds, we log-transform time in order to remove task dependancies on the scale of temporal propagation, to account for stream non-linearities and to alleviate exclusion of temporal outliers.
3.6 Context-enhanced Encoding
Literature has shown the effectiveness of enhancing word-level representations hierarchically through context-level learnt dynamics (Zheng et al., 2020; Wu et al., 2021; Ng et al., 2023). To this effect we introduce a second-layer of stream-level position embeddings, $\mathbf{s}^{11}$, to re-instate the absolute sequence position of each post for context-enhanced modeling. These are fed into a global context-aware layer, essentially a word-level transformer layer. Since the [CLS] tokens of each post are stream-aware, they contextualise the token-level representations based on the temporal and global learnt dynamics, obtaining: $[\mathbf{H}^{12}_{1,i-w+1},...,\mathbf{H}^{12}_{1,i}]$. To fully model the stream dynamics given the now context-enhanced [CLS] tokens, we employ a last layer of Temporal Rotary MHA resulting in $[\mathbf{H}^{{}^{\prime}12}_{1,i-w+1},...,\mathbf{H}^{{}^{\prime}12}_{1,i}]$. Lastly, we adapt the Gated Context Fusion (Gate&Norm) mechanism by Zheng et al. (2020) to fuse both the utterance word-level informed ($\mathbf{H}^{12}_{CLS}$) and the stream utterance-level informed ($\mathbf{H}^{{}^{\prime}12}_{CLS}$) [CLS] tokens through element-wise multiplication $\odot$:
$\mathbf{g}=\sigma(W_{g}[\mathbf{H}^{12}_{CLS};\mathbf{H}^{{}^{\prime}12}_{CLS}])$
$\mathbf{C}^{G}_{CLS}=\textrm{LayerNorm}[(1-\mathbf{g})\odot{H}^{12}_{CLS}+\mathbf{g}\odot{H}^{{}^{\prime}12}_{CLS}]$
3.7 Network Fine-Tuning
Although the proposed architecture can in principle be applied to any Transformer-based model, we select BERT (Devlin et al., 2018) as the foundation model and initialise all word-level weights. Literature on longitudinal context-aware classification has shown the importance of efficiently combining the current utterance with historical information (Sawhney et al., 2020, 2021a; Tseriotou et al., 2023). We thus concatenate the local stream-agnostic [CLS] token of the current utterance from the 10-th layer, $\mathbf{C}^{L}_{CLS}$, (obtained through typical BERT Pooling) with the obtained global stream-enhanced [CLS], $\mathbf{C}^{G}_{CLS}$ (Fig. 2). This final representation is fed through two fully connected layers with ReLU activation and dropout (Srivastava et al., 2014). The architecture is fine-tuned for each classification task (§4) using alpha-weighted focal loss (Lin et al., 2017), to assign more importance to minority classes and alleviate class imbalance.
4 Experiments
4.1 Tasks and Datasets
We test our model on three different longitudinal tasks of different temporal granularity: 1) Stance Switch Detection – identification of switches in overall user stance around a social media claim, 2) Moments of Change (MoC) – identification of mood changes through users’ online posts and 3) Conversation Topic Shift – conversation diversion identification. We adopt a real-time prediction formulation (see §3.1) to assess system ability to perform early change detection in real-world scenarios. Table 1 provides the detailed statistics of each dataset showing the different degree of temporal granularity and dataset specifics. More details are provided in Appendix A.
Stance Switch Detection: Introduced by Tseriotou et al. (2024) this binary classification task tracks the support of posts towards a rumourous claim to identify switches in the overall stance. We use the LRS dataset based on RumourEval-2017 (Gorrell et al., 2019) and use N-Sw: to denote absence of switch and Sw: (switch) to denote a shift in the number of opposing vs supporting posts.
Moments of Change (MoC): Introduced by Tsakalidis et al. (2022b) this 3-class classification task assesses the mood changes of users given their online posting timelines. We use the TalkLife dataset with: IS: (switches) sudden mood shift from positive to negative (or v.v.); IE: (escalations) gradual mood progression from neutral or positive/negative to more positive/negative; and O: no mood change. In both datasets each post is accompanied by its timestamp.
Conversation Topic Shift: Given a corpus of conversations between humans, this binary classification task identifies utterance relevance to the main conversation topic versus another (derailment). We use the Topic Shift-MI (Mixed-Initiative) dataset (Konigari et al., 2021) annotated on a subset of Switch-board (Godfrey et al., 1992; Calhoun et al., 2010). This has a single but different major topic for each conversation, with M: (major) utterances belonging to the main topic and R: (rest) utterances that pertain to a different topic. Here conversations are not timestamped.
Dataset LRS TalkLife Topic Shift MI # Data Points 5,568 18,604 12,536 # Timelines 325 500 74 Mean (median) 17.1 37.2 169.4 Timeline Length in Posts (13) (30) (153.5) Mean (median) 1h 26m 40s 6h 51m 11s - Time inbetween Posts 1m 39s 59m 38s (-) Mean (median) 6.5 IS:1.8, IE:4.0 60.5 # Minority Events/Timeline (0) (IS:1, IE:1) (51.5)
Table 1: Statistics of Datasets.
4.2 Baselines and Experimental Setup
We select baselines that are both post-level and stream-level (current post and recent history window, see §3.1).
Post-level: Random: post classification based on class distribution. BERT/RoBERTa: BERT (Devlin et al., 2018) or RoBERTa (Liu et al., 2019) fine-tuned using focal loss (Lin et al., 2017).
Llama2-7B-U (5/10-shot): Llama2-7B-chat-hf LLM (Touvron et al., 2023) with a crafted prompt for each dataset and few-shot (5/10) examples based on the distribution of input text, randomly sampled from each label distribution of each dataset, following Min et al. (2022). MistralInst2-7B-U(5/10-shot): MistralInst2-7B-U (Jiang et al., 2023): LLM with few shot examples. Same setting as for Llama2-7B-chat.
Stream-level: FFN History: Feed-forward on the concatenation of SBERT (Reimers and Gurevych, 2019) embeddings of the current post and the average of the window stream. SWNU (Tseriotou et al., 2023): feeds convolution-informed expanding path signature streams of SBERT representations and time into a BiLSTM modeling information progression. Seq-Sig-Net (Tseriotou et al., 2023): Sequential Network of SWNU modeled through a BiLSTM for long-term dependencies. BiLSTM: with a single layer on window stream SBERT embeddings. Llama2-7B-S (5-shot): Few-shot Llama2-7B-chat-hf prompted with the recent history of window 5 provided with each shot for context. MistralInst2-7B-S (5/10-shot): Same context-based setting as for Llama2-7B-S using Mistral-7B-Instruct-v0.2.
Evaluation: In line with published literature we report F1 scores for model performance, per class and macro-averaged. For each dataset we perform 5-fold cross validation with train/dev/test sets consisting of different timelines. We run and report the performance of each model on the exact same four random seeds (0,1,12,123) and report the average result on the test set. Appendix D provides information about implementation details and hyperparameter search.
5 Results and Discussion
5.1 Comparison against baselines
We present results for TempoFormer and baselines in Table 2. TempoFormer is the most performant in all three tasks based on macro-averaged F1. We note that recurrent models based on pre-trained BERT representations (BiLSTM for LRS and Topic Shift MI and Seq-Sig-Net for TalkLife), ranked second best. These models have been the SOTA for these datasets (Tseriotou et al., 2023, 2024). While the datasets are of different sizes, temporal characteristics, timeline length and change event distribution (see Table 1), TempoFormer retains its high performance, showcasing its generalisability for real-time change detection. Importantly, our model has the highest F1 for all minority classes, with the exception of Topic Shift MI, where other baselines have higher class-specific F1 scores for M but much lower F1 for R. Since TempoFormer operates on a contextual window of recent posts we select the appropriate window for each stream based on a window analysis, reported in §5.2.
We distinguish baselines into post and stream-level ones, noticing that smaller fine-tuned Language Models, even as simple as an FFN, allowing for stream-level context, score consistently better than post-level ones - with the exception of RoBERTa for TalkLife. This consistent finding underscores the importance of developing contextually informed representations for change detection. Few-shot prompted LLMs have consistently lower performance than smaller fine-tuned LMs, in line with reported poor temporal abilities of LLMs (Jain et al., 2023; Bian et al., 2023). For post-level, while Mistral’s performance improved from 5 to 10-shot, it is still barely above the random baseline and significantly behind BERT and RoBERTa. For LRS and Topic Shift MI the stream-level 5 and 10-shot Mistral performance increases, but falls way short of BERT/RoBERTa and all the stream-level models, indicating that although sequential context is important it is not modelled appropriately using LLMs. In line with (Wenzel and Jatowt, 2024), Llama2 suffers from generating responses outside the predefined classes, resulting in very low performance. TempoFormer demonstrates a generalisable architecture that enhances word-level post representations given the context, while modelling effectively the interplay between linguistic and temporal dynamics.
LRS TalkLife Topic Shift MI Model N-Sw Sw macro-avg IE IS O macro-avg M R macro-avg Post-level Random 61.4 37.5 49.5 11.2 4.5 84.4 33.4 35.9 63.9 49.9 Llama2-7B-U (5-shot) 22.4 50.6 36.5 10.1 7.5 31.9 16.5 46.6 45.4 46.0 MistralInst2-7B-U (5-shot) 71.4 28.0 49.7 23.3 4.1 67.8 31.7 46.4 44.6 45.5 Llama2-7B-U (10-shot) 8.8 52.5 30.7 12.8 6.2 31.3 16.7 48.5 39.5 44.0 MistralInst2-7B-U (10-shot) 71.2 30.5 50.8 27.6 3.5 72.1 34.4 42.6 55.7 49.1 BERT 69.0 45.3 57.1 43.9 28.1 86.8 52.9 36.0 70.0 53.0 RoBERTa 68.2 46.4 57.3 46.3 30.4 86.6 54.4 34.5 70.2 52.4 Stream-level FFN History 71.6 52.8 62.2 45.4 27.1 88.0 53.5 39.4 70.1 54.8 SWNU 75.5 55.5 65.5 48.0 29.3 89.5 55.6 38.7 66.0 52.3 Seq-Sig-Net 74.7 58.9 66.8 48.4 30.2 89.5 56.0 37.4 66.7 52.1 BiLSTM 75.0 60.7 67.8 46.1 27.0 89.2 54.1 37.8 73.8 55.8 Llama2-7B-S (5-shot) 2.2 50.2 26.2 15.5 7.6 24.2 15.7 52.6 1.3 27.0 MistralInst2-7B-S (5-shot) 58.3 50.2 54.3 22.0 4.6 70.0 32.2 42.3 57.3 49.8 MistralInst2-7B-S (10-shot) 54.4 51.8 53.1 23.4 3.5 74.9 33.9 37.8 63.7 50.8 TempoFormer (ours) 75.9 62.0 68.9 50.0 32.4 88.8 57.1 41.6 70.7 56.1
Table 2: (Best) F1-scores across all tasks. Stream-level models are applied on the optimal window, per dataset.
5.2 Window Length
Figure 3: TempoFormer Results for Different Contextual Window Sizes.
Since stream-based models operate on recent context, selecting appropriate contextual windows to include in the stream is important. Following Tseriotou et al. (2024) we determine window selection based both on model performance and dataset characteristics. Fig. 3 demonstrates TempoFormer’s F1 performance on windows of 5, 10 and 20 recent posts (see Table 11 for full results). While LRS and Topic Shift MI both benefit from the large window of 20 posts (blue) with clear performance gains overall and for the minority classes, TalkLife demonstrates better performance over a window of 10 (green). The optimal window findings for LRS and TalkLife are consistent with (Tseriotou et al., 2024). These differences are attributed to dataset characteristics (Table 1) and the mean number of change events in timelines, which need to be captured within the contextual windows. This analysis informs our stream-level experiments and at the same time demonstrates the flexibility of TempoFormer with respect to contextual window length. We recommend exploratory analysis according to dataset characteristics for appropriate window selection for new datasets.
5.3 Ablations Study
In Table 3 we present an ablation study to assess each of TempoFormer’s components.
Temporal Rotary Multi-head Attention (MHA): By using the vanilla sequential distance version of RoPE in Multi-head attention instead of the temporal one, for the timestamped datasets, we see a drop in performance. This showcases the advantage of modeling linguistic streams while accounting for their temporal dynamics and the success of temporally distant RoPE. The relatively small drop in performance is due to the secondary role of temporal dynamics compared to the linguistic evolution in change detection.
RoPE MHA: By further replacing the RoPE MHA with the vanilla version of MHA we see a significant drop in performance (in macro-avg): -2.9% for LRS, -1.2% for TalkLife and -0.6% for Topic Shift MI, demonstrating the success of RoPE on its own. We postulate that this signifies the ability of RoPE to enable MHA integration in architectures without the need for normalisation and FFN in a full transformer layer.
Stream embeddings: Removing these embeddings from TempoFormer layers brings noticeable performance drops in all datasets, showcasing the importance of propagating position information in building stream-aware and context-enhanced post embeddings. Topic Shift MI shows the largest drop of -1.4% among all its ablated models. Since this dataset does not obtain sequential signal from temporal dynamics, it relies on stream embeddings to model the distance between consecutive posts.
Gate&Norm operation updates the stream post-level [CLS] tokens with post word-level information, which is better informed by the word-level dynamics. This fuses together the word and stream dynamics in a gated learnable way. Large performance drops for all tasks when we ablate this component shows the importance of multi-level fusion.
LRS TalkLife Topic Shift MI Models N-Sw Sw macro-avg IE IS O macro-avg M R macro-avg TempoFormer 75.9 62.0 68.9 50.0 32.4 88.8 57.1 41.6 70.7 56.1 $\neg$¬Temporal RoPE 75.5 62.0 68.7 49.3 31.7 88.7 56.6 - - - $\neg$¬RoPE MHA 74.1 57.9 66.0 48.0 31.5 88.2 55.9 39.6 71.4 55.5 $\neg$¬Stream embed. 75.4 59.0 67.2 49.4 31.7 89.2 56.8 38.9 70.5 54.7 $\neg$¬Gate&Norm 74.5 61.3 67.9 49.8 31.1 88.7 56.6 40.7 69.6 55.2
Table 3: Ablation Studies for TempoFormer based on F1 with one component ablated at a time for all datasets.
5.4 The curious case of recurrence
Since longitudinal and change detection models have so far heavily relied on recurrence-based architectures, we evaluate the effect of recurrence on models jointly trained for stream and post-level representations. To do so we adapt RoBERT (Pappagari et al., 2019), originally developed for long document classification, which applies recurrence over BERT’s pooled [CLS] for each post. Here, we propose RoTempoFormer, a modification of RoBERT, that uses recurrence over TempoFormer’s pooled [CLS] for each post. Both RoBERT and RoTempoFormer are stream-level, recurrence-based models. We report results in Table 4.
RoTempoFormer consistently outperforms RoBERT for all datasets. RoTempoFormer strikes the right balance between jointly modeling context-aware post representations and recurrence in stream dynamics. Only for LRS do recurrence-based models have a better performance than TempoFormer. To examine this phenomenon we measure the diversity of each dataset with respect to its content and report it in Table 5. We report the BERTScore (Zhang et al., 2019b) and Cosine similarity between SBERT pairs of representations as well as the Outlier metric (Larson et al., 2019; Stasaski et al., 2020) on SBERT which measures the Euclidean distance between the (unseen) posts in the test set and the mean training corpus across folds and seeds for all datasets. Thus we assess both the semantic diversity and test set diversity. Across all metrics we consistently see that TalkLife is the most and LRS the least diverse. We postulate that for more diverse datasets like TalkLife, RoBERT has a really low performance, while it performs much better on less diverse ones. This could be due to: 1) overfitting due to recurrence and 2) inability of RoBERT to jointly model diverse context-aware representations, while capturing their evolution. RoTempoFormer, maintains its high performance, striking a good balance between modeling the context-aware post-level and the timeline-level dynamics. Importantly, we show that TempoFormer can be used as the foundation for temporal representation learning in other architectures.
LRS TalkLife Topic Shift MI model N-Sw Sw macro-avg IE IS O macro-avg M R macro-avg TempoFormer 75.9 62.0 68.9 50.0 32.4 88.8 57.1 41.6 70.7 56.1 RoBERT 75.8 62.3 69.0 36.7 3.3 88.4 42.8 33.3 75.7 54.5 RoTempoFormer 76.2 63.6 69.9 47.1 27.5 88.3 54.3 36.6 73.2 54.9
Table 4: Results (macro-avg F1) on Recurrent-based Language Models, including TempoFormer (non-recurrent) for comparison. Best scores are marked.
Dataset BERTScore $\downarrow$ Cosine Sim. $\downarrow$ Outlier $\uparrow$ LRS .457 .245 .867 TalkLife .358 .123 .934 Topic Shift MI .385 .188 .896
Table 5: Diversity Scores per Dataset.
5.5 Model Adaptability
model IE IS O macro-avg BERT 43.9 28.1 86.8 52.9 RoBERTa 46.3 30.4 86.6 54.4 TempoFormer (BERT) 50.0 32.4 88.8 57.1 TempoFormer (RoBERTa) 52.4 36.9 87.3 58.8
Table 6: Results (macro-avg F1) on TalkLife using BERT vs RoBERTa as the base model for TempoFormer.
To examine the flexibility of the TempoFormer stream-level and context-enhanced layers beyond the BERT architecture, we use TempoFormer with RoBERTa (roberta-base). Specifically, we allow the first 10 RoBERTa layers to model post-level (local) dynamics and modify its top two layers to capture stream dynamics. Since in Table 2, TalkLife benefits from the use of RoBERTa over BERT at the post-level, we examine if this gain also transfers to the TempoFormer. Summarising results in Table 6, we show that the RoBERTa-based TempoFormer achieves a new SOTA of 58.8% macro-avg F1, +1.7% over the BERT-based TempoFormer. This increase is in line with the +1.5% performance increase between vanilla BERT and RoBERTa macro-avg F1. Importantly, the increase in overall F1 is driven by clear performance gains in the IE and IS minority classes, further demonstrating the success and adaptable nature of TempoFormer in identifying changes over time.
6 Conclusion
We introduce TempoFormer, a transformer-based model for change detection operating on textual (timestamped) streams. Importantly we do so by avoiding recurrence, and only modifying the last two layers of the transformer. Furthermore, TempoFormer has the ability to model the temporal distance between textual units through a modification of rotary positional embeddings. The model achieves new SOTA, outperforming recurrent and LLM-based models on three different change detection tasks with datasets of varying temporal granularity and linguistic diversity, without loss in generalisability. We demonstrate its usability as a foundation model in other architectures, striking the right balance between word-level, post-level and stream-level linguistic and temporal dynamics. Lastly, we showcase its flexibility in terms of base model integration, further boosting stream-level performance on par with post-level gains.
Limitations
While TempoFormer shows SOTA performance on three different tasks and datasets of diverse temporal granularity involving change detection, namely: social media overall stance shift, user mood change detection and open conversation major topic shift detection, we are yet to evaluate its performance on a wider range of tasks and datasets. Additionally, although we demonstrate strong performance in datasets as small as  5,500 data points, we believe that our model, as most machine learning models, benefits from larger corpora in training where we can more meaningfully fine-tune the inter and intra-post relationships to model the dataset’s linguistic style and change intricacies. TempoFormer models post dynamics through a predefined stream window, identified through understanding the characteristics of a dataset via preliminary experiments. The need for initial exploration can be limiting compared to a dynamic window setting. Furthermore, despite the fact that our implementation is flexible and can be applied to different encoder architectures, the codebase is built in PyTorch, therefore imposing the constraint of PyTorch-only frameworks. On the classification front, we operate on a supervised setting therefore assuming the availability of annotated data which can be expensive to obtain especially from experts. Regarding evaluation, we focus on post-level metrics, and have not yet considered metrics more appropriate for longitudinal tasks and streams Tsakalidis et al. (2022b). Lastly, since our model operates by fine-tuning a pre-trained transformer-based model, like BERT, it automatically assumes the availability of such model in the language of the dataset/interest (English in our case), which might not be the case for low-resource languages.
Ethics Statement
The performance of our model, TempoFormer, is demonstrated on three datasets: LRS, TalkLife and Topic Shift MI. The LRS dataset is based on the publicly available RumourEval 2017 dataset (Gorrell et al., 2019) for stance detection, while the Topic Shift MI dataset is also a publicly available dataset based on human to human open domain conversations. Since the TalkLife dataset contains sensitive and personal user data, the appropriate Ethics approval was received from the Institutional Review Board (IRB), followed by data anonymisation and appropriate sensitive data sharing procedures. Access to this dataset was granted and approved by TalkLife  [^3] https://www.talklife.com/ through licensing for research purposes associated with the corresponding submitted proposal. All examples in the paper are paraphrased. Models were built on a secure server with authorised user-only access, while the labeled TalkLife dataset and the developed models are not indented to be made publicly available in order avoid any potential risks of unintended use.
Appendix A Dataset Specifics
Since we are following 5-fold cross validation the test set consists of  20% of the datapoints. For LRS and Topic Shift MI the remaining data are split 25%/75% between dev/train sets and for TalkLife they are split 33.3%/66.7% between dev/train sets. The difference between these percentages is in order to ensure that we have substantial training data for LRS and Topic Shift MI in each fold as these are relatively small datasets in size. Splitting between train/dev/test is stratified so that all timeline examples belong only to one of the sets, therefore the above percentages are approximate (not exact).
Appendix B Libraries
All experiments were ran under the same Python 3.10.12 environment including these libraries: pandas=1.5.2, matplotlib=3.7.1, pip=23.2.1, scikitlearn=1.2.0, pytorch=2.0.1, pytorch-cuda=11.8, transformers=4.35.0, tokenizers=0.14.1, huggingface-hub==0.20.3
For Seq-Sig-Net and SWNU baselines we used the Sig-Networks package and its environment as reported in Tseriotou et al. (2024).
Appendix C Computational Infrastructure
The experiments for the LRS and TalkLife datasets were ran on a machine with 2 NVIDIA A40 GPUs of 48GB GPU RAM each, 96 cores and 256 GB of RAM.
The experiments for the Topic Shift dataset were ran on machine with 3 NVIDIA A30 GPUs of 24GB GPU RAM each, 40 cores and 384 GB of RAM.
Appendix D Experimental Details
Implementation Details In our experiments for all models we train on 4 epochs with early stopping and patience 3, gradient accumulation and focal loss with $\gamma=2$ and alpha of $\sqrt{1/p_{t}}$ where $p_{t}$ is the probability of class $t$ in the training data Tseriotou et al. (2023). For Transformer-based models we use the AdamW optimiser (Loshchilov and Hutter, 2017) and a linear scheduler and for the rest we use the Adam optimiser (Kingma and Ba, 2014). The models are implemented using Pytorch (Paszke et al., 2019).
For TempoFormer we use bert-base-uncased. We build our custom model with Huggingface’s (Wolf et al., 2019) BERT classes and RoPE Llama classes (Touvron et al., 2023) as a starting point. All applicable BERT defaults are kept unchanged, using max length of 512 and 12 attention heads. For the classification feedforward-network we use two 64-dimensional layers and a dropout of 0.1 with ReLU. Following an initial space search, learning rate is selected using grid-search on: $[1e^{-5},5e^{-6}]$.
BERT/RoBERTa: Fine-tuned versions of bert-base-uncased/roberta-base using a grid search over learning rates of $\in$ $[1e^{-6},5e^{-6},1e^{-5}]$.
FFN History: Following Tseriotou et al. (2024), we perform hyperparameter search over learning rates $\in$ $[1e^{-3},5e^{-4},1e^{-4}]$ and hidden dimensions $\in$ $[[64,64],[128,128],[256,256],[512,512]]$, over 100 epochs with a batch size of 64 and a dropout rate of 0.1.
SWNU and Seq-Sig-Net: We perform a hyperparameter search over: learning rates $\in$ $[0.0005,0.0003]$, feed-forward hidden dimensions of the two layers $\in$ $[[32,32],[128,128]]$, LSTM hidden dimensions of SWNU units $\in$ $[10,12]$, convolution-1d reduced dimensions $\in$ $[6,10]$ and BiLSTM hidden dimensions for Seq-Sig-Net of $\in$ $[300,400]$. Models were developed using the log-signature, time encoding in the path as well as concatenated at its output for LRS and TalkLife and sequence index in the path for Topic Shift MI. We use 100 epochs with a batch size of 64 and a dropout rate of 0.1.
BiLSTM: Following Tseriotou et al. (2024), we perform hyperparameter search over learning rates $\in$ $[1e^{-3},5e^{-4},1e^{-4}]$ and hidden dimensions $[200,300,400]$, over 100 epochs with a batch size of 64 and a dropout rate of 0.1.
SBERT: SentenceBERT (SBERT) representations were used for different baselines (Reimers and Gurevych, 2019) in order to obtain semantically meaningful post-level embeddings. We use 384-dimensional embeddings through all-MiniLM-L6-v2 from the sentence_transformers library.
RoBERT: Following Pappagari et al. (2019) we develop RoBERT with the exact same parameters as in the original paper and a grid search through learning rates $\in$ $[1e^{-6},5e^{-6},1e^{-5}]$. We follow the same grid search for RoTempoFormer.
Appendix E LLM Prompts
To construct Mistral classification prompts we follow the recommended classification prompts as per provided guidelines  [^4] https://docs.mistral.ai/guides/prompting_capabilities/#classification. For constructing the Llama prompts we experimented with multiple prompts per dataset and identified the ones with the most stable performance. For fairer performance assessment we apply post-processing in LLM predictions to bucket them in the corresponding classification class (e.g. if the LLM generates esc we mark it as an escalation). In Tables 7, 8, 9, 10 we provide our LLM prompts for the LRS dataset.
Table 7: MistralInst2-7B-U for n-shot Post/Utterance-level prompting
MistralInst2-7B-U Template
You are a helpful, respectful and honest assistant for labeling online Twitter conversations between users. Given the online post of a user in a conversation stream around a rumourous claim on a newsworthy event which it is discussed by tweets in the stream, determine if in the current post there is a switch with respect to the overall stance. Answer with "none" for either the absence of a switch or cases where the numbers of supporting and opposing posts are equal and with "switch" for switch between the total number of oppositions (querying or denying) and supports or vice versa. Your task is to assess and categorize post input after <<<>>>much-less-thanmuch-greater-than<<<>>>< < < > > > into one of the following predefined outputs:
none
switch
You will only respond with the output. Do not include the word "Output". Do not provide explanations or notes.
Here are some examples:
Input: post example 1
Output: post label 1
$\cdots$
Input: post example n
Output: post label n
Table 8: MistralInst2-7B-S for n-shot Stream-level prompting
MistralInst2-7B-S Template
You are a helpful, respectful and honest assistant for labeling online Twitter conversations between users. Given the most recent online conversation history between users around a rumourous claim on a newsworthy event, determine if the most recent input user post is a switch with respect to the overall conversation stance. Answer with "none" for either the absence of a switch or cases where the numbers of supporting and opposing posts are equal and with "switch" for switch between the total number of oppositions (querying or denying) and supports or vice versa. Your task is to assess and categorize post input after <<<>>>much-less-thanmuch-greater-than<<<>>>< < < > > > into one of the following predefined outputs:
none
switch
You will only respond with the output. Do not include the word "Output". Do not provide explanations or notes.
Here are some examples:
Conversation History:
$u_{a-4}$
$u_{a-3}$
$u_{a-2}$
$u_{a-1}$
Input: post example 1, $u_{a}$
Output: post label 1
$\cdots$
Conversation History:
$u_{b-4}$
$u_{b-3}$
$u_{b-2}$
$u_{b-1}$
Input: post example n, $u_{b}$
Output: post label n
Table 9: Llama2-7B-U for n-shot Post/Utterance-level prompting
Llama2-7B-U Template
$<s>[INST]<<SYS>>$
You are a helpful, respectful and honest assistant for labeling online Twitter conversations between users.
$<</SYS>>$
Given the online post of a user in a conversation stream around a rumourous claim on a newsworthy event which it is discussed by tweets in the stream, determine if in the current post there is a switch with respect to the overall stance.
Answer with "none" for either the absence of a switch or cases where the numbers of supporting and opposing posts are equal and with "switch" for switch between the total number of oppositions (querying or denying) and supports or vice versa.
Example 1:
Input: post example 1
Output: post label 1
$\cdots$
Example n:
Input: post example n
Output: post label n
Only return "none" or "switch".
Limit the answer to 1 word.
$[/INST]$
$</s>$
Table 10: Llama2-7B-S for n-shot Stream-level prompting
Llama2-7B-S Template
$<s>[INST]<<SYS>>$
You are a helpful, respectful and honest assistant for labeling online Twitter conversations between users.
$<</SYS>>$
Given the most recent online conversation history between users around a rumourous claim on a newsworthy event, determine if the most recent input user post is a switch with respect to the overall conversation stance.
Answer with "none" for either the absence of a switch or cases where the numbers of supporting and opposing posts are equal and with "switch" for switch between the total number of oppositions (querying or denying) and supports or vice versa.
Example 1:
Conversation History:
$u_{a-4}$
$u_{a-3}$
$u_{a-2}$
$u_{a-1}$
Input: post example 1, $u_{a}$
Output: post label 1
$\cdots$
Example n:
Conversation History:
$u_{b-4}$
$u_{b-3}$
$u_{b-2}$
$u_{b-1}$
Input: post example n, $u_{b}$
Output: post label n
Only return "none" or "switch".
Limit the answer to 1 word.
$[/INST]$
$</s>$
Appendix F Window Results
Full results for the window analysis are presented in Table 11.
LRS TalkLife Topic Shift window N-Sw Sw avg IE IS O avg M R avg 5 69.9 55.0 62.5 50.1 27.9 88.7 55.6 38.3 71.1 54.7 10 73.0 56.4 64.7 50.0 32.4 88.8 57.1 40.3 71.1 55.7 20 75.9 62.0 68.9 49.5 32.0 88.8 56.8 41.6 70.7 56.1
Table 11: F1 scores for TempoFormer on all datasets for different window sizes. Best scores are marked.